I need to write the equation x^15=8x^5 as a polynomial equation, in the form p(x)=0. I don't understand how to do this and need help. Thanks!
Answer by nerdybill(7384)

If you started with:
x^15=8x^5
simply move all the terms to one side by subtracting 8x^5 from both sides:
x^15 - 8x^5 = 0 (this is then your polynomial as a function of x)
.
You could "factor" the left side to give you:
x^5(x^10 - 8) = 0